Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
-12x + 27y = 60
-6x – 6y = 30

Answer by mananth(16946)

-12 x + 27 y = 60 .............1

-6 x -6 y = 30 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 9
-24 x 54 y = 120
-54 x -54 y = 270
Add the two equations
-78 x = 390
/ -78
x = -5
plug value of x in (1)
-12 x + 27 y = 60
60 + 27 y = 60
27 y = 60 -60
27 y = 0
y = 0
(-5,0) is the solution